OpenStudy (anonymous):

help: completing the square: 2/3x^2-2x+7/3=0

OpenStudy (lalaly):

multiply each side by 3 \[\large{2x^2-6x+7}=0\]now divide by two\[\large{x^2-3x+\frac{7}{2}}=0\]\[\large{(x-\frac{3}{2})^2-(\frac{3}{2})^2}+\frac{7}{2}=0\]\[\large{(x-\frac{3}{2})^2+\frac{5}{4}=0}\]

OpenStudy (anonymous):

is that already finish?

OpenStudy (anonymous):

do you need to find the value of x now?

OpenStudy (anonymous):

completing the square has been done by lalaly

OpenStudy (anonymous):

yes jimmyrep

OpenStudy (anonymous):

ok - i'm not very good at equation thing so (x - 3/2)^2 = -5/4 take square root of both sides: x - 3/2 = +- sqrt (-5/4) x = +- sqrt (-5/4) + 3/2 roots are 3/2 + (sqrt5) i/2 or 3/2 - (sqrt5) i/2

OpenStudy (anonymous):

sorry - i was trying to use the equation function but messed up ! are you ok with the answer i gave?

OpenStudy (anonymous):

sorry, but i cant totally understand it

OpenStudy (anonymous):

did you understand it as far as the fourth line?

OpenStudy (anonymous):

no

OpenStudy (lalaly):

\[\large{(x-\frac{3}{2})^2}=\frac{-5}{4}\]

OpenStudy (lalaly):

\[\large{x-\frac{3}{2}}=\sqrt{\frac{-5}{4}}\]

OpenStudy (lalaly):

\[x=\frac{3}{2} \pm \sqrt{\frac{-5}{4}}\]

OpenStudy (anonymous):

are you ok with lalaly's last 3 posts?

OpenStudy (lalaly):

\[\underline{\large{Hope so}}\]